At t= 0, a block of mass m is projected with speed v on a rough long horizontal plank (of mass 3m) placed on smooth ground, as shown. Work done by friction on the block between t= 0 and t= infinity equals
1
Expert's answer
2019-03-21T10:54:26-0400

Using momentum conservation law:


"p_{initial}=p_{final} \rightarrow
mv=(m+3m)v_f"

"v_f=\frac{mv}{4m}=0.25v"

where, v_f - speed of two blocks at t=infinity


Total work done by friction force equals to the loss in kinetic energy:


"A=\Delta E=\frac{mv^2}{2}-\frac{(m+3m)v_f^2}{2}"

"A=\frac{mv^2}{2}-\frac{mv^2}{8}"

"A=\frac{3mv^2}{8}"

So, total work done by friction force is

"A=3mv^2\/8"